Competitive downhill ski racers are subjected to severe wind chill, as are average skiers and snowboarders to a lesser degree. Assuming a downhill racer is going 96.6 kmph (60 mph), which is coasting on some runs, and the air temperature is −18°C (0°F),

What is their time to experience frostbite given these conditions? ______________ minutes.
1
Expert's answer
2020-11-04T14:13:31-0500

A. 96.6km/h * 1,000 / 3,600 = 26.4 m/s

a = V^2 / 2s

a = (26.4 m/s)^2 / 2 * 105 = 3.31 m/s^2

F = ma

800 kg * 3.31 m/s^2

2655 N


B. a = v^2 / 2s

a = (96.6 * 5 / 18)^2 / 2(2)

a = 174.09 m/s^2

F = ma

F = (800 kg)(174.09)

F = 139274.7

force in (b) / force in (a) = 139274.7 N / 2655 N

52.45